For which values of k, is the function f, defined as below, continuous at x = 2 ?
F(x) = 3-kx, 1 is less than equal to x is less than 2.
(X^2/4) -3 ,x is greater than equal to 2
Further, at which other points in [1,∞[ is f continuous, and why?
1
Expert's answer
2019-08-19T08:26:43-0400

"3-kx" and "\frac{x^2}{4}-3" are continuous all over their domains because these are polynomials, thus "F" is continuous all over "[1;2[\cup]2;\infty[" for an arbitrary "k".

So we need to find "\lim\limits_{x\to 2-0}3-kx" and "\lim\limits_{x\to 2+0}\left(\frac{x^2}{4}-3\right)" and set them equal to achieve "F" to be continuous at "x=2" . So

"\lim\limits_{x\to 2-0}3-kx=3-2k"

"\lim\limits_{x\to 2+0}\left(\frac{x^2}{4}-3\right)=\frac{2^2}{4}-3=1-3=-2" , so

"3-2k=-2"

"k=\frac52"